OpenStudy (anonymous):

x-3=12 then whats the value of x-3=3?

OpenStudy (kira_yamato):

12?

OpenStudy (anonymous):

is this a real question

OpenStudy (saifoo.khan):

x = 12 + 3 x = 15. Now, x = 15 into equation 2. 15-3 = 3 \[12 \ne 3\]

OpenStudy (anonymous):

what? do you mean and thank you!